OpenStudy (anonymous):

integral of x+1/x^2+1

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ x+1 }{ x^2 + 1}\] This one?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Separate it into two, try it from here. \[\int\limits_{}^{} \frac{ x }{ x^2+1} + \int\limits_{}^{} \frac{ 1 }{ x^2+1}\]

OpenStudy (anonymous):

i got up to this part and i know what to do for 1/x^2+1 but when i do x/x^2+1 i get confused

OpenStudy (anonymous):

Use substitution

OpenStudy (anonymous):

did that and i got u=x^2+1 and du=2xdx which is du/2=xdx

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

does it end up being 1/2 the integral of 1/u du?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank you i was looking online but yahoo answers was just confusing me

OpenStudy (anonymous):

np

hartnn (hartnn):

thats the difference between yahoo answers (answering site) and OpenStudy(study site). we make sure you learn and can do similar problems on your own by interacting. Good work! @ECE :) \(Welcome ~to ~OpenStudy!!~~\ddot \smile \) @arrowhead94

OpenStudy (anonymous):

thank you!!! @hartnn